OpenStudy (anonymous):

The Formula for degrees in Celsius is: C = 5/9 (f-32) Where F stands for degrees Fahrenheit. Please help me learn the steps to solve for f and Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (anonymous):

you need to manipulate the equation to get the formula for degrees in Fahrenheit. so: from: C = 5/9(F-32) (times both sides by the inverse of 5/9 (which is 9/5), this will cancel out the right 5/9) C x 9/5 = F - 32 Now just add 32 to both sides to get rid of the -32 on the right and you will have: C x 9/5 + 32 = F Now just plug in 20 degrees into the C for Celsius and solve for Fahrenheit :)

OpenStudy (anonymous):

Wow. Thank you very much. That was extremely helpful.

OpenStudy (anonymous):

No problem at all :)